Price and cost MC Market quantity

Part 1

(a) Which area represents consumer surplus under perfect competition? ABH? ACG? DCG?

(b) Which area represents producer surplus under perfect competition? DCG? BHJD? ACG?

Part 2  

(a) Which area represents consumer surplus under monopoly? BHJD? ACG? ABH?

(b) Which area represents producer surplus under monopoly? BHJD? CDG? ABH?

Part 3  

Which area represents the deadweight loss associated with a monopoly? HGJ? BHGC? HGFE?
Answer #1

P-1:

In a perfectly competitive market, the equilibrium is (P = MC)

Qa) Consumer surplus (CS) is ACG.

CS is the triangular area bounded by the maximum price on demand curve and equilibrium price; the maximum price is A on the demand curve D; the equilibrium price is C, since the corresponding G is on the (P = MC)) equilibrium.

Qb) Producer surplus (PS) is DCG

PS is the area bounded by the suppliers’ minimum price and equilibrium price (corresponding to P = MC) on MC curve.

P-2:

In a monopoly market, the equilibrium is (MR = MC)

Qa) Consumer surplus (CS) is ABH.

CS is the triangular area bounded by the maximum price on demand curve and equilibrium price; the maximum price is A on the demand curve D; the equilibrium price is B, since the corresponding H is on the (MC = MR) equilibrium.

Qb) Producer surplus (PS) is BHJD

PS is the area bounded by the suppliers’ minimum price and equilibrium price (corresponding to MC = MR) on MC curve.

P-3:

Total surplus in perfect competition = CS + PS = ACG + DCG = ADG

Total surplus in monopoly = CS + PS = ABH + BHJD = AHJD

Deadweight loss = Total surplus in perfect competition - Total surplus in monopoly

                            = ADG – AHJD

                            = HGJ (Answer)